# Incommensurate Twisted Bilayer Graphene: emerging quasi-periodicity and stability

### v1

@@ VI.2 Feynman rules for the renormalized expansion
Some clusters are single vertices (either $\nu$ or $\tau$) and are called trivial clusters. The clusters that contain internal lines are called non-trivial clusters. As per the construction above, if $h^{ext}_{T}$ is the largest of the scales of the external lines of a non-trivial cluster $T$, all its internal lines have a scale $h>h^{ext}_{T}$; $h_{T}$ is the largest scale of the propagators internal to the cluster $T$. A non-trivial cluster $T$ contains sub-clusters ${\widetilde{T}}\subset T$. We call a cluster ${\widetilde{T}}\subset T$ a maximal cluster if there is no other cluster $\bar{T}$ such that ${\widetilde{T}}\subset\bar{T}\subset T$.
@@
${\cal R}W^{h}({\bf k}+p_{F,i}^{\omega})=k^{2}\int_{0}^{1}\partial^{2}W(t{\bf k})$ (55)
@@ VI.3 Convergence of the series and persistence of the cones
In order to bound the contribution of a Feynman graph we note that $|\bar{g}^{(h)}(k)|\leq C\gamma^{-h}$ (by (50)); therefore the product of propagators in the Feynman graph is bounded by
@@
where $M_{T}$ is the number of maximal non-resonant clusters contained in $T$, $R_{T}$ is the number of maximal resonant clusters contained in $T$, and $\prod_{T\mathrm{\ n.t.}}$ is a product over non-trivial clusters. Note that the trivial clusters do not contribute, as they contain no internal lines. The effect of the ${\cal R}:=\mathds{1}$-$\mathcal{L}$ (recall (42)) operation on the non-trivial resonant clusters produces an extra $\gamma^{2(h_{T}^{\text{ext}}-h_{T})}$ (the gain $h_{T}^{\mathrm{ext}}$ comes from the extra $(k-p_{F,i}^{\omega})^{2}$ terms (which are on the scale of the external lines) and the $h_{T}$ loss from the extra second derivative (which are on the scale of the internal lines)). Therefore, we get an extra factor:
@@
    If $i_{1}=i_{2}=1$ and $k_{1},k_{2}$ are the momenta of the external lines; assume that $k_{1}$ is in the first Brillouin zone and $k_{1}=:\bar{k}_{1}+p_{F,1}^{\omega_{1}}$. Moreover $k_{2}:=k_{1}+\widehat{m}_{T}b^{\prime}+lb=:\bar{k}_{2}+p_{F,1}^{\omega_{2}}$, with $l$ chosen such that $k_{2}$ is in the first Brillouin zone; then, if $\widehat{m}_{T}\not=0$, by (28),
@@
    $|p_{F,1}^{\omega_{1}}-p_{F,1}^{\omega_{2}}+\widehat{m}_{T}b^{\prime}+lb|\geq{C_{0}\over|\widehat{m}_{T}|^{\tau}}$ (72)
@@
    for some constant $c_{1}>0$. On the other hand if $\widehat{m}_{T}=0$ but $\omega_{1}\not=\omega_{2}$ then $l=0$ so $\gamma^{h^{\mathrm{ext}}_{T}}\geq\frac{1}{2}|p_{F,1}^{\omega_{1}}-p_{F,1}^{\omega_{2}}|=\frac{2\pi}{3\sqrt{3}}$ (recalling (10)). Thus, this eventuality does not occur provided $\gamma$ is large enough.
@@
    In the case $i_{1}=1$ and $i_{2}=2$ and $k_{1},k_{2}$ are the momenta of the external lines; assume that $k_{1}$ is in the first Brillouin zone and $k_{1}=:\bar{k}_{1}+p_{F,1}^{\omega_{1}}$. Moreover $k_{2}:=k_{1}+\bar{l}b+mb^{\prime}=:\bar{k}_{2}+p_{F,2}^{\omega_{2}}$, with $m$ chosen in such a way that $k_{2}$ is in the first Brillouin zone; then, if $\bar{l}\not=0$, by (28),
@@
    $|p_{1,F}^{\omega_{1}}-p_{2,F}^{\omega_{2}}+\widehat{l}_{T}b+mb^{\prime}|\geq{C_{0}\over|\widehat{l}_{T}|^{\tau}}$
@@
for some constant $c_{2}$. Therefore,
@@
In addition, $\sum_{T\,\text{n.t.}}M_{T}\leq q$, since the clusters are nested in each other and for two clusters to be different they must differ by at least one vertex. Now, let us introduce $M_{T}^{\tau}$ as the number of maximal non-resonant trivial clusters (i.e. maximal $\tau$-vertices) contained in $T$, and use the trivial bound $3M_{T}\leq 2M_{T}+M_{T}^{\tau}$ along with (82) to obtain
@@
The crucial point is that the sum over the scales $h$ can be performed summing over all the differences, taking into account that the scale $h$ is fixed. Finally the sum over the $l,m$ is done using the factor $L(\underline{l},\underline{m})]^{1\over 2}$. (The gain term $\gamma^{h_{T}M_{T}^{\tau}}$ is dropped, as it does not lead to any significant gain.) In conclusion the bound on a graph with $q$ vertices is $C_{4}^{q}\gamma^{h}|\lambda|^{q}$ assuming that $|\nu_{h}|,|Z_{h}-1|,|v_{h}-1|,|w_{h}-1|\leq C|\lambda|$.
VI.4 Beta function and Schwinger functions
We are left with checking our assumption on $Z_{h},\nu_{h},w_{h},v_{h}$. We know that $v_{i,\omega,h}=v_{i,\omega,h+1}-iz_{h,\omega,a,b,i,1}$ with $z_{h,\omega,a,b,i,1}$ expressed by the sum of renormalized Feynman graphs $\Gamma$ such that the maximal scale of the clusters is $h+1$, an extra derivative is applied (which costs a factor $\gamma^{-h}$) and the momenta of the external lines is fixed equal to $p_{F}^{\omega}$. Moreover by the compact support of the propagator there is at least a $\tau$ vertex, as the $k=0$ value of a graph wih only $\nu$ vertice is zero; therefore the analogue of (86) becomes
@@ VI.4 Beta function and Schwinger functions
$\gamma^{h\over 2}\sum^{*}_{\underline{h},\atop\underline{l},\underline{m}}[L]^{1\over 2}|\lambda|^{q}(CC_{3})^{q}\prod_{T\,\text{n.t.}}\gamma^{{1\over 2}(h_{T}^{\text{ext}}-h_{T})}.$ (89)
Estimating the sum as above, we find that $|z_{h,\omega,a,b,i,1}|\leq C_{5}\lambda\gamma^{h\over 2}$, and $v_{i,\omega,h}=v_{i,\omega,0}-i\sum_{h^{\prime}}z_{h^{\prime},\omega,a,b,i,1}$ hence $v_{i,\omega,h}=v_{i,\omega,0}+O(\lambda)$; moreover the limiting value is reached exponentially fast $v_{i,\omega,h}=v_{i,\omega,-\infty}+O(\lambda\gamma^{h/2})$. A similar argument holds for $Z_{h},w_{h}$. Not
It remain to discuss the flow of $\nu_{h}$; we can write, see (46),
$W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,p_{F,i}^{\omega})=\gamma^{h+1}\nu_{h+1}+{\widetilde{W}}_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,p_{F,i}^{\omega})$ (90)
@@
with $\beta^{h}_{\nu}=\gamma^{-h}{\widetilde{W}}_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,p_{F,i}^{\omega})$ is given by the sum with $q\geq 2$ of terms bounded by (89). We have to prove that it is possible to choose the counterterms $\nu_{\omega,\alpha,\alpha^{\prime},i}$ so that $\nu_{h,\omega,\alpha,\alpha^{\prime},i}$ is bounded by $C\lambda$ for any scale $h$. Indeed from (91) we get, $h\leq-1$
@@
$\widehat{S}_{j,j}(\mathbf{k}+\mathbf{p}_{F,j}^{\omega}))=\sum_{h=-\infty}^{0}[\widehat{g}^{(h)}((\mathbf{k}+\mathbf{p}_{F,j}^{\omega}))+r^{h}(\mathbf{k})$ (94)
where $r^{h}(\mathbf{k})$ includes the contribution of term withs at least a vertex. We can replace in $g^{(h)}((\mathbf{k}+\mathbf{p}_{F,j}^{\omega}))$ the $v_{i,\omega,h},w_{i,\omega,h},Z_{i,\omega,h}$ with $v_{i,\omega,-\infty},w_{i,\omega,-\infty},Z_{i,\omega,-\infty}$ obtaining the dominant term in (36); the subdominant term is obtained both from the term containing the difference betwen $v_{i,\omega,h}-v_{i,\omega,-\infty}$, $z_{i,\omega,h}-w_{i,\omega,-\infty}$, $Z_{i,\omega,h}-Z_{i,\omega,-\infty}$, which have an extra factor $O(\lambda\gamma^{h/2})$, or the terms with at least a vertex which have at least a $\nu$ or a non resonant trivial vertex, with an extra $O(\gamma^{h/2})$ from the bounds after (87).
VII Conclusion
Theoretical analyses of TBG are based on the assumption of the stability of the Weyl semimetallic phase, leading to the formulation of continuum effective models. However in lattice TBG models wth generic angles there is an emerging quasi-periodicity manifesting themself in large momenta Umklapp interactions that almost connect the Dirac points, similar to the ones appearing in electronic systems with quasi-periodic potential. Such terms are neglected in the continuum semimetallic approximations.
In this paper we have rigorously established the stability of the semimetallic phase in a lattice model, taking into full account the large momenta Umklapp interactions. The analysis is based on number theoretical properties of irrationals combined with a Renormalization Group analysis, and requires that the interlayer hopping is weak and short ranged and the the angles are chosen in a large measure set. The effect of the interaction is to produce a finite renormalization of the Dirac points and velocities. Non perturbative effects are excluded as the series are shown to be convergent. Compared to the Aubry-André or similar models, the number theoretical analysis is much more involved due to the peculiar structure of the small divisors.
The stability of the Weyl phases provides a justification of the use of continnum models under the above assumptions. In addition, the present analysis paves the way to a more accurate evaluation of the velocities as functions of the angles, talking into account lattice or higher orders effects, and the effect of many body interactions, whose interplay with the emerging quasi-periodicity could lead to interesting phases.
@@ Appendix A Fourier transform of the interlayer hopping
Noting that $\widehat{c}_{2,k_{1}+lb-mb^{\prime},\alpha}\equiv\widehat{c}_{2,k_{1}+lb,\alpha}$, $\widehat{c}_{1,k_{2}^{\prime}+mb^{\prime}-lb,\alpha}\equiv\widehat{c}_{1,k_{2}^{\prime}+mb^{\prime},\alpha}$ we finally obtain (II).
@@
Appendix B Proof of Lemma IV.1
To prove lemma IV.1, we will first prove a general result on a Diophantine condition for a generic function from $0,2\pi)$ to $\mathbb{R}^{2}$. We will then apply this result to $|p_{F,i}^{\omega},p_{F,j}^{\omega^{\prime}}+lb+mb^{\prime}|$, viewed as a function of $\theta$, for the various values of $i,j,\omega,\omega^{\prime}$.
@@ B.1 Diophantine condition from $\mathbb{R}$ to $\mathbb{R}^{2}$
$\zeta:=\mathbb{Z}^{2}\setminus(\{0\}\cup\mathcal{C}_{\xi}({\textstyle\frac{\pi}{4}}))$ (102)
(the reason why we choose $\Omega$ in this way will become apparent in the proof of Lemma B.1 below).
@@
$R_{1}:=\frac{C_{3}+\frac{|x^{\prime}(\theta)|}{|f^{\prime}(\theta)|}}{\cos(\alpha+\frac{\pi}{4})}$ (105)
@@
$\min_{0<|k|<R_{2}}|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|-\eta|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}})>0$ (106)
@@
$R_{2}:=\eta+\frac{|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|}{|{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}})}$ (107)
@@
$|[\theta_{0},\theta_{1}]\setminus\mathcal{D}(x,f)|\leq O(C_{3})+O({\textstyle\frac{C_{1}}{C_{3}\beta}})$ (108)
where the constants in $O(\cdot)$ depend only on $\theta_{0}$, $\theta_{1}$, $x$, $f$, $\alpha$, $\epsilon$, $\tau$, $\eta$. In particular, if we choose $C_{3}\ll\theta_{1}-\theta_{0}$ and $C_{1}\ll(\theta_{1}-\theta_{0})^{2}$, then $\mathcal{D}(x,f)$ fills most of $\theta_{0},\theta_{1}]$.
@@
$g_{l,k}(\theta)=|x(\theta)+l+k\cdot f(\theta)|$ (109)
@@
$|\mathcal{D}_{\Omega}^{c}(x,f)|=\sum_{k,l}^{*}\int_{-C_{1}|k|^{-\tau}}^{C_{1}|k|^{-\tau}}\frac{1}{g_{l,k}^{\prime}}dg_{l,k}$ (110)
where $\sum^{*}_{k,l}$ has the constraint that $\exists\theta\in[\theta_{0},\theta_{1}]$ such that $g_{k,l}(\theta)\in[-C_{1}|k|^{-\tau},C_{1}|k|^{-\tau}]$. Therefore
@@
In addition, the number of values of $l$ such that $g_{k,l}(\theta)\in[-C_{1}|k|^{-\tau},C_{1}|k|^{-\tau}]$ is bounded by $C_{2}|k|$ for some constant $C_{2}$ (which depends only on $\theta_{0},\theta_{1},x,f$), and so
@@
To do so, $\Omega$ must be chosen appropriately: we wish for $k\cdot f^{\prime}$ to stay as far away from $x^{\prime}$ as possible. Now, it cannot avoid it entirely, as $k\cdot f^{\prime}$ will cover all possible values as $k$ varies in $\mathbb{Z}^{2}\setminus\{0\}$. By choosing $\Omega$ as in (101), we ensure that $k\cdot f^{\prime}$ may only approach $x^{\prime}$ for large values of $k$. In doing so, we can estimate $\mathcal{D}_{\Omega}^{c}$: we split the sum over $\mathbb{Z}^{2}\setminus\{0\}$ into a sum over $\zeta$ and a sum over its complement $\zeta^{c}\equiv\mathbb{Z}^{2}\cap\mathcal{C}_{\xi}(\frac{\pi}{4})$, and compute a bound for each case.
@@
If, on the other hand, $k\in\zeta^{c}\equiv\mathbb{Z}^{2}\cap\mathcal{C}_{\xi}(\frac{\pi}{4})$,
$|k\cdot f^{\prime}(\theta)|\geq|k||f^{\prime}(\theta)|\cos(\alpha+{\textstyle\frac{\pi}{4}})$ (114)
@@
$|x^{\prime}(\theta)+k\cdot f^{\prime}(\theta)|\geq|k||f^{\prime}(\theta)|\cos(\alpha+{\textstyle\frac{\pi}{4}})-|x^{\prime}(\theta)|.$ (115)
@@
$|k|\geq\frac{C_{3}+\frac{|x^{\prime}(\theta)|}{|f^{\prime}(\theta)|}}{\cos(\alpha+\frac{\pi}{4})}\equiv R_{1}$ (116)

### v2

@@ VI.2 Feynman rules for the renormalized expansion
Some clusters are single vertices (either $\nu$ or $\tau$) and are called trivial clusters. The clusters that contain internal lines are called non-trivial clusters. As per the construction above, if $h^{ext}_{T}$ is the largest of the scales of the external lines of a non-trivial cluster $T$, all its internal lines have a scale $h>h^{ext}_{T}$; $h_{T}$ is the largest scale of the propagators internal to the cluster $T$. A non-trivial cluster $T$ contains sub-clusters ${\widetilde{T}}\subset T$. We call a cluster ${\widetilde{T}}\subset T$ a maximal cluster if there is no other cluster $\bar{T}$ such that ${\widetilde{T}}\subset\bar{T}\subset T$. See e.g. [24], [25], [28] for more details.
@@
${\cal R}\widehat{W}^{h}({\bf k}+\mathbf{K}^{i}_{\omega})=k^{2}\int_{0}^{1}\partial^{2}\widehat{W}(t{\bf k})$ (58)
@@ VI.3 Convergence of the series and persistence of the cones
In this section we want to prove the convergence of the series, extending methods previously used [24], [25], [28] for fermions in a quasi-periodic potential. In order to bound the contribution of a Feynman graph we note that $|\bar{g}^{(h)}(k)|\leq C\gamma^{-h}$ (by (53)); therefore the product of propagators in the Feynman graph is bounded by
@@
where $M_{T}$ is the number of maximal non-resonant clusters contained in $T$, $R_{T}$ is the number of maximal resonant clusters contained in $T$, and $\prod_{T\mathrm{\ n.t.}}$ is a product over non-trivial clusters. Note that the trivial clusters do not contribute, as they contain no internal lines. The effect of the ${\cal R}:=\mathds{1}$-$\mathcal{L}$ (recall (45)) operation on the non-trivial resonant clusters produces an extra $\gamma^{2(h_{T}^{\text{ext}}-h_{T})}$ (the gain $h_{T}^{\mathrm{ext}}$ comes from the extra $(k-K^{i}_{\omega})^{2}$ terms (which are on the scale of the external lines) and the $h_{T}$ loss from the extra second derivative (which are on the scale of the internal lines)). Therefore, we get an extra factor:
@@
    If $i_{1}=i_{2}=1$ and $k_{1},k_{2}$ are the momenta of the external lines; assume that $k_{1}$ is in the first Brillouin zone and $k_{1}=:\bar{k}_{1}+K^{1}_{\omega_{1}}$. Moreover $k_{2}:=k_{1}+\widehat{m}_{T}b^{\prime}+lb=:\bar{k}_{2}+K^{1}_{\omega_{2}}$, with $l$ chosen such that $k_{2}$ is in the first Brillouin zone; then, if $\widehat{m}_{T}\not=0$, by (4),
@@
    $|K^{1}_{\omega_{1}}-K^{1}_{\omega_{2}}+\widehat{m}_{T}b^{\prime}+lb|\geq{C_{0}\over|\widehat{m}_{T}|^{\tau}}$ (75)
@@
    for some constant $c_{1}>0$.
    On the other hand if $\widehat{m}_{T}=0$ but $\omega_{1}\not=\omega_{2}$ then $l=0$ so $\gamma^{h^{\mathrm{ext}}_{T}}\geq\frac{1}{2}|K^{1}_{\omega_{1}}-K^{1}_{\omega_{2}}|=\frac{2\pi}{3\sqrt{3}}$ (recalling (13)). Thus, this eventuality does not occur provided $\gamma$ is large enough.
@@
    In the case $i_{1}=1$ and $i_{2}=2$ and $k_{1},k_{2}$ are the momenta of the external lines; assume that $k_{1}$ is in the first Brillouin zone and $k_{1}=:\bar{k}_{1}+K^{1}_{\omega_{1}}$. Moreover $k_{2}:=k_{1}+\bar{l}b+mb^{\prime}=:\bar{k}_{2}+K^{2}_{\omega_{2}}$, with $m$ chosen in such a way that $k_{2}$ is in the first Brillouin zone; then, if $\bar{l}\not=0$, by (4),
@@
    $|K^{1}_{\omega_{1}}-K^{2}_{\omega_{2}}+\widehat{l}_{T}b+mb^{\prime}|\geq{C_{0}\over|\widehat{l}_{T}|^{\tau}}$
@@
for some constant $c_{2}$; note that such constant is proportional, up to some numerical constant, to $C_{0}^{\tau}$, as follows from (76). Therefore,
@@
In addition, $\sum_{T\,\text{n.t.}}M_{T}\leq q$, since the clusters are nested in each other and for two clusters to be different they must differ by at least one vertex.
Now, let us introduce $M_{T}^{\tau}$ as the number of maximal non-resonant trivial clusters (i.e. maximal $\tau$-vertices) contained in $T$, and use the trivial bound $3M_{T}\leq 2M_{T}+M_{T}^{\tau}$ along with (85) to obtain
@@
The crucial point is that the sum over the scales $h$ can be performed summing over all the differences, taking into account that the scale $h$ is fixed. Finally the sum over the $l,m$ is done using the factor $L(\underline{l},\underline{m})]^{1\over 2}$. (The gain term $\gamma^{h_{T}M_{T}^{\tau}}$ is dropped, as it does not lead to any significant gain.) In conclusion the bound on a graph with $q$ vertices is $C_{4}^{q}\gamma^{h}|\lambda|^{q}$ assuming that $|\nu_{h}|,|Z_{h}-1|,|v_{h}-1|,|w_{h}-1|\leq C|\lambda|$. The constant $C_{4}$ is proportional to $C_{0}^{4\tau}$ as one can check collecting the estimates above.
VI.4 Beta function and imaginary time Green functions
We are left with checking our assumption on $Z_{h},\nu_{h},w_{h},v_{h}$. We know that $v_{i,\omega,h}=v_{i,\omega,h+1}-iz_{h,\omega,a,b,i,1}$ with $z_{h,\omega,a,b,i,1}$ expressed by the sum of renormalized Feynman graphs $\Gamma$ such that the maximal scale of the clusters is $h+1$, an extra derivative is applied (which costs a factor $\gamma^{-h}$) and the momenta of the external lines is fixed equal to $K_{\omega}$. Moreover by the compact support of the propagator there is at least a $\tau$ vertex, as the $k=0$ value of a graph with only $\nu$ vertices is zero; therefore the analogue of (89) becomes
@@ VI.4 Beta function and imaginary time Green functions
$\gamma^{h\over 2}\sum^{*}_{\underline{h},\atop\underline{l},\underline{m}}[L]^{1\over 2}|\lambda|^{q}(CC_{3})^{q}\prod_{T\,\text{n.t.}}\gamma^{{1\over 2}(h_{T}^{\text{ext}}-h_{T})}r.$ (92)
Estimating the sum as above, we find that
$|z_{h,\omega,a,b,i,1}|\leq C_{5}\lambda\gamma^{h\over 2}\quad v_{i,\omega,h}=v_{i,\omega,0}-i\sum_{h^{\prime}}z_{h^{\prime},\omega,a,b,i,1}$ (93)
[…]
It remains to discuss the flow of $\nu_{h}$; we can write, see (49),
$W_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,K^{i}_{\omega})=\gamma^{h+1}\nu_{h+1}+{\widetilde{W}}_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,K^{i}_{\omega})$ (94)
@@
with $\beta^{h}_{\nu}=\gamma^{-h}{\widetilde{W}}_{i,i,0,\alpha,\alpha^{\prime}}^{h,\omega,\omega}(0,K^{i}_{\omega})$ is given by the sum with $q\geq 2$ of terms bounded by (92). We have to prove that it is possible to choose the counterterms $\nu_{\omega,\alpha,\alpha^{\prime},i}$ so that $\nu_{h,\omega,\alpha,\alpha^{\prime},i}$ is bounded by $C\lambda$ for any scale $h$. Indeed from (95) we get, $h\leq-1$
@@
$\widehat{S}_{j,j}(\mathbf{k}+K^{j}_{\omega}))=\sum_{h=-\infty}^{0}\widehat{g}^{(h)}(\mathbf{k}+\mathbf{K}^{j}_{\omega})+r^{h}(\mathbf{k})$ (98)
where $r^{h}(\mathbf{k})$ includes the contribution of term with at least a vertex. We can replace in $g^{(h)}((\mathbf{k}+\mathbf{K}^{j}_{\omega}))$ the $v_{i,\omega,h},w_{i,\omega,h},Z_{i,\omega,h}$ with $v_{i,\omega,-\infty},w_{i,\omega,-\infty},Z_{i,\omega,-\infty}$ obtaining the dominant term in (33); the subdominant term is obtained both from the term containing the difference between $v_{i,\omega,h}-v_{i,\omega,-\infty}$, $z_{i,\omega,h}-w_{i,\omega,-\infty}$, $Z_{i,\omega,h}-Z_{i,\omega,-\infty}$, which have an extra factor $O(\lambda\gamma^{h/2})$, or the terms with at least a vertex which have at least a $\nu$ or a non resonant trivial vertex, with an extra $O(\gamma^{h/2})$ from the bounds after (90).
VII Conclusions
We have considered the role of large-momentum-transfer Umklapp terms. These terms, which nearly connect the Fermi points of different layers, are typically neglected in effective continuum descriptions but could destroy the Dirac cones. We rigorously find that, for finite interlayer coupling, the semimetallic phase is stable provided the angles belong to a fractal set of large relative measure (decreasing with the hopping) characterized by the validity of a Diophantine condition. Our result therefore ensures that for weak hopping for most (in the sense specified above) angles the semimetallic behaviour persists even in presence of the large momentum Umklapp terms which almost connect the Dirac points. The result provides a partial justification to the effective continuum description of TBG in which terms are neglected.
It would be of course a very interesting problem to consider hopping outside the regime we are considering to see if there can be some transition to a different phase due to such terms, as it happens for fermions in a 1D quasi-periodic Aubry potential; or to detect in the weak coupling regime some sign of fractality following from the above number theoretical considerations. In addition, the present analysis paves the way to a more accurate evaluation of the velocities as functions of the angles, taking into account lattice or higher-order effects, and the effect of many-body interactions, whose interplay with the emerging quasi-periodicity could lead to interesting phases.
@@ Appendix A Fourier transform of the interlayer hopping
Noting that $\widehat{c}_{2,k_{1}+lb-mb^{\prime},\alpha}\equiv\widehat{c}_{2,k_{1}+lb,\alpha}$, $\widehat{c}_{1,k_{2}^{\prime}+mb^{\prime}-lb,\alpha}\equiv\widehat{c}_{1,k_{2}^{\prime}+mb^{\prime},\alpha}$ we finally obtain (16).
@@
Appendix B Proof of Lemma V.1
To prove lemma V.1, we will first prove a general result on a Diophantine condition for a generic function from $0,2\pi)$ to $\mathbb{R}^{2}$. We will then apply this result to $|K^{i}_{\omega},K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|$, viewed as a function of $\theta$, for the various values of $i,j,\omega,\omega^{\prime}$.
@@ B.1 Diophantine condition from $\mathbb{R}$ to $\mathbb{R}^{2}$
$\zeta:=\mathbb{Z}^{2}\setminus(\{0\}\cup\mathcal{C}_{\xi}(\beta))$ (106)
where $\beta\in(\alpha,\frac{\pi}{2}-\alpha)$ (the reason why we choose $\Omega$ in this way will become apparent in the proof of Lemma B.1 below).
@@
$R_{1}:=\frac{C_{3}+\frac{|x^{\prime}(\theta)|}{|f^{\prime}(\theta)|}}{\cos(\alpha+\beta)}$ (109)
@@
$\min_{0<|k|<R_{2}}|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|-\eta|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{2}}-\beta)>0$ (110)
@@
$R_{2}:=\eta+\frac{|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|}{|{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+\frac{\pi}{2}-\beta)}$ (111)
@@
$|[\theta_{0},\theta_{1}]\setminus\mathcal{D}(x,f)|\leq O(C_{3})+O({\textstyle\frac{C_{1}}{C_{3}}})$ (112)
where the constants in $O(\cdot)$ depend only on $\theta_{0}$, $\theta_{1}$, $x$, $f$, $\alpha$, $\beta$, $\epsilon$, $\tau$, $\eta$. In particular, if we choose $C_{3}\ll\theta_{1}-\theta_{0}$ and $C_{1}\ll(\theta_{1}-\theta_{0})^{2}$, then $\mathcal{D}(x,f)$ fills most of $\theta_{0},\theta_{1}]$. If we further choose $C_{3}\sim\sqrt{C_{1}}\ll(\theta_{1}-\theta_{0})$, then the measure of the complement of $\mathcal{D}(x,f)$ is $O(\sqrt{C_{1}})\ll\theta_{1}-\theta_{0}$.
@@
$g_{l,k}(\theta)=x(\theta)+l+k\cdot f(\theta)$ (113)
@@
$|\mathcal{D}_{\Omega}^{c}(x,f)|\leq\sum_{k,l}^{*}\int_{-C_{1}|k|^{-\tau}}^{C_{1}|k|^{-\tau}}\frac{1}{g_{l,k}^{\prime}}dg_{l,k}$ (114)
where $\sum^{*}_{k,l}$ has the constraint that $\exists\theta\in[\theta_{0},\theta_{1}]$ such that $g_{l,k}(\theta)\in[-C_{1}|k|^{-\tau},C_{1}|k|^{-\tau}]$. Therefore
@@
In addition, the number of values of $l$ such that $g_{l,k}(\theta)\in[-C_{1}|k|^{-\tau},C_{1}|k|^{-\tau}]$ is bounded by $C_{2}|k|$ for some constant $C_{2}$ (which depends only on $\theta_{0},\theta_{1},x,f$), and so
@@
To do so, $\Omega$ must be chosen appropriately: we wish for $k\cdot f^{\prime}$ to stay as far away from $x^{\prime}$ as possible. Now, it cannot avoid it entirely, as $k\cdot f^{\prime}$ will cover all possible values as $k$ varies in $\mathbb{Z}^{2}\setminus\{0\}$. By choosing $\Omega$ as in (105), we ensure that $k\cdot f^{\prime}$ may only approach $x^{\prime}$ for large values of $k$. In doing so, we can estimate $\mathcal{D}_{\Omega}^{c}$: we split the sum over $\mathbb{Z}^{2}\setminus\{0\}$ into a sum over $\zeta$ and a sum over its complement $\zeta^{c}\equiv\mathbb{Z}^{2}\cap\mathcal{C}_{\xi}(\beta)$, and compute a bound for each case.
@@
If, on the other hand, $k\in\zeta^{c}\equiv\mathbb{Z}^{2}\cap\mathcal{C}_{\xi}(\beta)$,
$|k\cdot f^{\prime}(\theta)|\geq|k||f^{\prime}(\theta)|\cos(\alpha+\beta)$ (118)
@@
$|x^{\prime}(\theta)+k\cdot f^{\prime}(\theta)|\geq|k||f^{\prime}(\theta)|\cos(\alpha+\beta)-|x^{\prime}(\theta)|.$ (119)
@@
$|k|\geq\frac{C_{3}+\frac{|x^{\prime}(\theta)|}{|f^{\prime}(\theta)|}}{\cos(\alpha+\beta)}\equiv R_{1}$ (120)